I need help on finding a slope-intercept formula with the slope of 2 and the point that it passes through being (3,2). Can you please help me?
Answer by kennedysj(5)

Slope-intercept form is y=mx+b
If your slope is 2 first substitute this into the form for the m so you get
y = 2x+b
now you know the values of x and y from the point you are given (3,2)
We can substitute 3 for x and 2 for y so we have
2 = 2(3) + b.
We now need to find what the value of b is so we solve
2= 2(3)+ b or
2 = 6 + b or
-4 = b.
we can substitute this value of b into our equation to get
y = 2x -4